# Conservation-informed Graph Learning for Spatiotemporal Dynamics Prediction

### v1

@@ 5.3. Evaluation metrics
where $\widehat{\mathbf{U}}\in\mathbb{R}^{N\times d\times T}$ and $\mathbf{U}\in\mathbb{R}^{N\times d\times T}$ denote the predicted and true system states, respectively; $cov[\cdot,\cdot]$ the covariance function; $\sigma_{\mathbf{u}}$ and $\sigma_{\hat{\mathbf{u}}}$ the standard deviations of $\mathbf{U}$ and $\widehat{\mathbf{U}}$, respectively; and $T$ the total number of considered time steps. In addition, introducing noise to the training data can stabilize the model and improve the generalization ability for learning spatiotemporal dynamics (Sanchez-Gonzalez et al., 2020; Pfaff et al., 2021). Therefore, we add a small amount of Gaussian noise in the training data to improve the CiGNN’s performance.
5.4. Training settings and computational resources
For a fair comparison, we train each model ten times and select the one with the median performance according to the evaluation of the validation datasets. We train all models for 1,000 epochs and set early stopping with a patience of 100 epochs. We utilize the GELU activation function (Paszke et al., 2019) for all MLP hidden layers , while linear activation is considered for the output layers. Additionally, layer normalization is applied after each MLP to improve the training convergence, except for those in the Decoder. All models are trained using the Adam Optimizer (Kingma and Ba, 2015) and the ReduceLROnPlateau learning scheduler (Paszke et al., 2019) which decays the learning rate by a factor of $0.8$. The encoder module consists of a 2-layer MLP with a hidden size of 128, and the decoder has a 2-layer MLP with a hidden size of 128. We only use 4 SymMPNN blocks for the processor, where 2 layer MLPs with a hidden size of 128 are employed as the node, edge and flux update functions. The same hyperparameters are used in MGN, MP-PDE, and CiGNN. Notably, all the hyperparameters are selected from various ranges. For example, the learning rate is selected from the set $[10^{-1},10^{-2},10^{-3},10^{-4},10^{-5}]$. More details about the hyperparameter test are given in Appendix Tables S4 to S8. The models were trained on the NVIDIA A100 GPU (80GB) on an Intel(R) Xeon(R) Platinum 8380 CPU (2.30GHz, 64 cores) server. More details are given in Appendix Section D.
5.5. Results
5.5.1. 2D Convection-Diffusion system
The results listed in Table 1 show that CiGNN consistently outperforms other baseline models in the context of prediction accuracy. Figure 3a displays the distribution of prediction errors as well as the predicted snapshots at a typical time step (0.995 s) by different methods. The CiGNN model lifts the prediction accuracy by 1$\sim$2 orders of magnitude in terms of MSE metric compared with all other baselines. Due to the meshless and point-wise nature of DeepONet, it falls short in scalable modeling of higher dimensional spatiotemporal dynamics. Given the squared domain with structured grid mesh, FNO produces a moderate prediction result in the situation with limited training data. The performance of MP-PDE aligns with that of MGN. Note that MP-PDE is a derivative of MGN with the primary distinction in their decoder components, e.g., MGN employs multilayer perceptron (MLP) while MP-PDE utilizes convolution for decoding. Despite discrepant performance of other models, CiGNN generalizes well to ICs and maintains a high prediction accuracy, which demonstrates notable strength and shows clear superiority over the baselines.
5.5.2. 3D Reaction-Diffusion system
5.5.3. 2D Flow past a cylinder
Figure 3. Error distributions and the system state snapshots predicted by CiGNN and other baselines. More results are shown in Appendix Figures S2S3. a, The 2D viscous Burgers equation example. b, The 3D Gray-Scott equation example. c, Schematic diagram of the CF example setup with the fluid flowing in from the left (inflow) to the right (outflow).
@@ 5.5. Results
5.5.4. 2D Hydrological dynamics forecast of the Black Sea
Table 2. Ablation results on four datasets. Here, the symbol “–” represents that Task #$\#2$ on CF dataset does not exist since multi-mesh strategy is not utilized. The bold values and underlined values represent the optimal and sub-optimal results on various datasets.
@@ 5.6. Ablation study
To gain deeper insights into how each component of the model affects its overall performance, we conducted an ablation study based on the aforementioned datasets. The results show that the removal of each module leads to deteriorated performance of the model as shown in Table 2. We summarize the observations as follows:
@@
We also consider two variants of CiGNN, namely, CiGNN- and CiGNN, where only the asymmetric information passing is considered in the SymMPNN blocks. As depicted in Appendix Table S10, the result shows the efficacy of the symmetric information-passing component. In the case of BS dataset, CiGNN outperforms CiGNN, which indicates that the effective learning of the inequality term in real-world scenarios is not fully guaranteed.
Moreover, we provide the ablation results on comparison of various temporal integration strategies shown in Table 3. It is evident that a 4-layer CiGNN model with the simplest 1D convolution filter-based strategy (denoted by C3, with minimal parameter overhead) achieves the best performance, surpassing other strategies including the 2D convolution filtering, TCN and LSTM, demonstrating that our temporal strategy (1D Conv filter) plays a key role in CiGNN to improve the model’s long-term rollout prediction accuracy. More detailed discussion can found in Appendix Subsection E.3.
[…]
Prediction results of CiGNN on real-world BS hydrological dataset.
@@
Case No. of layers Temporal strategy No. of parameters RMSE
@@ Acknowledgements.
The work is supported by the National Natural Science Foundation of China (No. 62276269, No. 92270118), the Beijing Natural Science Foundation (No. 1232009), the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0620103), and the National Key R&D Program of China (No. 2021ZD0110400). H. Sun would like to thank the support from the Fundamental Research Funds for the Central Universities and the Research Funds of Renmin University of China (No. 202230265). Y. Liu would like to thank the support from the Fundamental Research Funds for the Central Universities (No. E2EG2202X2). P. Ren would like to disclose that he was involved in this work when he was at Northeastern University, who has not been supported by Huawei Technologies.
Appendix A Background
The study of spatiotemporal dynamics aims to explore and elucidate the dynamic behavior, pattern formation, instability, and predictive capabilities of complex systems. Understanding spatiotemporal dynamics is crucial for addressing practical challenges in various domains, including physics, chemistry, engineering, climate science, etc. Recently, machine learning (ML) has shown great promise in modeling domain-specific scientific problem governed by partial differential equations (PDEs). In this paper, we focus on graph neural networks (GNNs) to learn spatiotemporal PDE systems.
A.1. Conventional GNNs
A classic GNN (Welling and Kipf, 2016) is composed of three main components: an embedding layer for the input features, a stack of GNN layers, and a final task-based layer. In a graph $G=(V,E)$ where $V$ contains $n=|V|$ nodes and $E$ is the set of edges, the connectivity of the graph is represented by the adjacent matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$. Here, $\mathbf{A}_{ij}=1$ if there exists an edge between the nodes $i$ and $j$, otherwise $\mathbf{A}_{ij}=0$. The degree matrix is denoted as $\mathbf{D}\in\mathbb{R}^{n\times n}$.
A.2. Conventional MPNNs
For simplicity, we describe the process of message-passing neural networks (MPNNs) (Gilmer et al., 2017) on the undirected graph $G$ with node features $\mathbf{h}_{i}$ and edge features $\mathbf{e}_{ij}$. The forward pass has two phases, a message-passing phase and an update phase. The message-passing phase is expressed by message function $\phi_{e}$ and the update phase is described by vertex update function $\phi_{v}$. The network layers are indexed by $l$ and $\mathbf{h}^{l=0}$ denotes the node features of the initial input layer. During these two phase, the hidden state $\mathbf{h}_{i}^{l+1}$ at node $i$ in the graph are updated by the messages $\mathbf{e}_{ij}^{l,*}$ and its last hidden state $\mathbf{h}_{i}^{l}$ according to
(S1a) $\mathbf{e}_{ij}^{l,*}$ $=\phi_{e}^{l}\left(\mathbf{h}_{i}^{l}\parallel\mathbf{h}_{j}^{l}\parallel\mathbf{e}_{ij}^{l}\right),$
(S1b) $\mathbf{h}_{i}^{l+1}$ $=\phi_{v}^{l}\left(\mathbf{h}_{i}^{l}\parallel\sum_{j\in\mathcal{N}_{i}}\mathbf{e}_{ij}^{l,*}\right),$
[…]
A.3. Variables Notation
Moreover, we present a summary of the variable notations used in our paper, as detailed in Table S1.
Table S1. Variables notation used in our paper.
Application Domain Variable Description Symbol Role
2D Burgers Equation x-component of velocity $u(x,y,t)$ Input / Predicted
y-component of velocity $v(x,y,t)$ Input / Predicted
2D Navier-Stokes Equation vorticity $w(x,y,t)$ Input/Predicted
2D Gray-Scott Equation concentrations of chemical species $U$ $u(x,y,t)$ Input/Predicted
concentrations of chemical species $V$ $v(x,y,t)$ Input/Predicted
3D Gray-Scott Equation concentrations of chemical species $U$ $u(x,y,z,t)$ Input/Predicted
concentrations of chemical species $V$ $v(x,y,z,t)$ Input/Predicted
2D Cylinder Flow Dataset x-component of velocity $u(x,y,t)$ Input/Predicted
y-component of velocity $v(x,y,t)$ Input/Predicted
pressure $p(x,y,t)$ Input / Predicted
2D Black Sea Dataset x-component of water flow velocity $u(x,y,t)$ Input/Predicted
y-component of water flow velocity $v(x,y,t)$ Input/Predicted
temperature in water depth of 12.5 meter ${T}(x,y,t)$ Input/Predicted
Global Field x-direction of space coordinate $x$ Input
y-direction of space coordinate $y$ Input
z-direction of space coordinate $z$ Input
time coordinate $t$ Input
time increment $\Delta t$ Input
space increment $\Delta\mathbf{x}$ Input
discrete timestamp at $k$th step $t_{k}$ Input
Appendix B Methodology
This section provides the supplementary information for the methodology described in the Main Text.
B.1. Preliminary of Graph Calculus
Let $G=(V,E)$ be a connected finite graph, which includes $|V|$ nodes and $|E|$ edges containing all edges between vertices in $V$. In the scalar or vector field, the sum of degrees is formulated as $\sum_{{u}_{i}\in{V}}deg({u}_{i})=2|E|$, where the symbol $deg({u}_{i})$ denotes the number of edges that are incident to the vertex ${u}_{i}$. Under the aforementioned conditions, we furnish the relevant definitions below.
Definition 1: With a vertex function f: $V\rightarrow\mathbb{C}$, the symbol $\nabla_{ij}f$ for ${j}\in\mathcal{N}_{i}$ is described as follows:
(S2) $\nabla_{ij}f=f\left({u}_{j}\right)-f\left({u}_{i}\right),$
where $\mathcal{N}_{i}$ denotes all other nodes within the neighborhood of node ${u}_{i}$ and we simplify the subscript ${u}_{i}{u}_{j}$ as $ij$.
Definition 2: With a vertex function f: $V\rightarrow\mathbb{C}$, the symbol $\nabla_{ij}f^{2}$ for ${j}\in\mathcal{N}_{i}$ is defined as
(S3) $\nabla_{ij}f^{2}=f\left({u}_{j}\right)^{2}-f\left({u}_{i}\right)^{2}.$
Definition 3: With any vertex function f: $V\rightarrow\mathbb{C}$ and g: $V\rightarrow\mathbb{C}$, the symbol $\nabla_{ij}\left(fg\right)$ for ${j}\in\mathcal{N}_{i}$ is formulated as follows:
(S4) $\nabla_{ij}\left(fg\right)$ $=f\left({u}_{j}\right)g\left({u}_{j}\right)-f\left({u}_{i}\right)g\left({u}_{i}\right)$
$=f\left({u}_{j}\right)\left(g\left({u}_{j}\right)-g\left({u}_{i}\right)\right)+\left(f\left({u}_{j}\right)-f\left({u}_{i}\right)\right)g\left({u}_{i}\right)$
$=f\left({u}_{j}\right)\nabla_{ij}\left(g\right)+\nabla_{ij}\left(f\right)g\left({u}_{i}\right)$
$=\left(\nabla_{ij}\left(f\right)+f\left({u}_{i}\right)\right)\nabla_{ij}\left(g\right)+\nabla_{ij}\left(f\right)g\left({u}_{i}\right)$
$=\nabla_{ij}\left(f\right)\nabla_{ij}\left(g\right)+f\left({u}_{i}\right)\nabla_{ij}\left(g\right)+\nabla_{ij}\left(f\right)g\left({u}_{i}\right).$
Definition 4: For ${j}\in\mathcal{N}_{i}$, the gradient of vertex function $f$ at any ${u}_{i}\in V$ is given by
(S5) $\nabla f\left({u}_{i}\right)=\left(\nabla_{ij}f\right).$
Definition 5: For ${{j}\in\mathcal{N}_{i}}$ on $G$, the scalar product of any vector field $\mathbf{W}\left({u}_{i}\right)=\left(\overline{w}_{ij}\right)$ and $\mathbf{U}\left({u}_{i}\right)=\left(\overline{u}_{ij}\right)$ is given by
(S6) $\mathbf{W}\left({u}_{i}\right)\cdot\mathbf{U}\left({u}_{i}\right)=\sum_{{j}\in\mathcal{N}_{i}}\overline{w}_{ij}\overline{u}_{ij}.$
Specifically, we have
(S7) $|\mathbf{W}\left({u}_{i}\right)|^{2}=\sum_{{j}\in\mathcal{N}_{i}}\overline{w}_{ij}^{2}.$
Definition 6: If $f$ is a scalar function (i.e., a real-valued or complex-valued function) and $\mathbf{W}\left({u}_{i}\right)$ is a vector field, their pointwise product can be written as follows:
(S8) $\left(f\mathbf{W}\right)\left({u}_{i}\right)=f\left({u}_{i}\right)\cdot\mathbf{W}\left({u}_{i}\right),$
where the function $f$ scales each component of the vector field $\mathbf{W}\left({u}_{i}\right)$ at every point ${u}_{i}$ in space.
Definition 7: For ${u}_{i}\in V$, the directional derivative of function $f$ along a vector field $\mathbf{W}\left({u}_{i}\right)$ is defined by
(S9) $\nabla_{\mathbf{W}}f\left({u}_{i}\right)=\mathbf{W}\cdot\nabla f\left({u}_{i}\right)=\sum_{{j}\in\mathcal{N}_{i}}\overline{w}_{ij}\nabla_{ij}f.$
Then, the directional derivative of a function $f$ along $\nabla f$ has the form of
(S10) $\nabla_{\nabla f}f\left({u}_{i}\right)=\sum_{{j}\in\mathcal{N}_{i}}\nabla_{ij}f\nabla_{ij}f=|\nabla f\left({u}_{i}\right)|^{2}.$
Definition 8: For the vector field $\mathbf{W}\left({u}_{i}\right)$ at ${u}_{i}\in V$, its divergence can be defined by
(S11) $div\left(\mathbf{W}\left({u}_{i}\right)\right)=\nabla\cdot\mathbf{W}\left({u}_{i}\right)=\sum_{{j}\in\mathcal{N}_{i}}\overline{w}_{ij}.$
Then, the divergence of $\nabla f\left({u}_{i}\right)$ at ${u}_{i}\in V$, also named the laplacian operator, is given by
(S12) $div\left(\nabla f\left({u}_{i}\right)\right)=\sum_{{j}\in\mathcal{N}_{i}}\nabla_{ij}f=\nabla^{2}f\left({u}_{i}\right).$
Divergence theorem: For ${j}\in\mathcal{N}_{i}$, ${u}_{i}\in V$, if $\overline{w}_{ij}=-\overline{w}_{ji}$ , we have
[…]
Proof.
Set $\overline{w}_{ij}=0$ for any ${u}_{j}\not\in\mathcal{N}\left({u}_{i}\right)$, then we have
(S14) $\sum_{{u}_{i}\in V}\sum_{{j}\in\mathcal{N}_{i}}\overline{w}_{ij}$ $=\sum_{{u}_{i}\in V}\sum_{{u}_{j}\in V}\overline{w}_{ij}$
$=\frac{1}{2}\sum_{{u}_{i}\in V}\sum_{{u}_{j}\in V}\overline{w}_{ij}+\frac{1}{2}\sum_{{u}_{i}\in V}\sum_{{u}_{j}\in V}\overline{w}_{ji}$
$=0.$
B.2. Conservative Property of CiGNN
@@
Theorem 1: If there is a continuously differentiable tensor field $\mathbf{F}$ on a neighborhood of $\mathcal{V}$, then we have that a volume integral over the volume $\mathcal{V}$ is equal to the surface integral over the boundary of the volume $\mathcal{V}$:
@@
where $flux_{i}$ denotes the total flow quantity through the $i$th node, and $flow_{i\cdot}$ and $flow_{\cdot i}$ the flows that leave and enter the node $i$. Subsequently, an easy corollary of Theorem 1 is described as follows.
Corollary 1: If the volume ${\mathcal{V}}$ can be partitioned into some separate parts $\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{n}$, the flux out $\Phi(\mathcal{V})$ of the original volume ${\mathcal{V}}$ should be equal to the sum of flux through each component volume, namely,
@@
where the edge weight function $w_{ij}$ associates to every edge ${e}_{ij}$ a value, namely, $w({u}_{i},{u}_{j})$. Similarly, $w_{ji}$ associates to every edge ${e}_{ji}$ a value, namely, $w({u}_{j},{u}_{i})$.
Definition 2: Based on the weighted graph derivative, we define the linear weighted gradient operator on graphs $\nabla_{w}:\mathcal{H}(V)\rightarrow\mathcal{H}(E)$ as:
@@
where $\mathcal{N}_{i}$ denotes the neighborhood node’s index set for node $i$. For simplicity, we simplified $F({u}_{i},{u}_{j})$ as $F_{ij}$ and $F({u}_{j},{u}_{i})$ is simplified by $F_{ji}$.
@@
where ${s}(\cdot)$ represents the source term. Then, substituting Eqs. S20 and S22 into Eq. S23, we get the following equation:
@@
B.3. Spatial and Temporal Strategies
The motivations of our network design are two-fold in terms of feature extraction of spatial patterns and temporal evolution, respectively.
B.3.1. Spatial strategy
Firstly, GNNs have shown great potential for modeling spatiotemporal systems thanks to their inherent capacity to capture the dynamics in the physical world using graph structures. However, a critical observation reveals that current GNN models tend to overemphasize node features while neglecting the crucial aspect of edge features (Hamilton et al., 2017). In particular, leveraging GNNs for modeling spatiotemporal dynamics should be able to include both explicit temporal changes in node states and implicit interactions via edges. The existing GNN models focus on processing edge features to align with data distribution without obeying a prior knowledge of physics principles. Hence, we introduce two forms of prior knowledge to ensure the GNN model conforming to the general conservation law, namely, symmetry (e.g., conservation) and asymmetry (e.g., entropy inequality), and design the SymMPNN block as a core graph learning module. These concepts draw inspiration from the well-known conservation law and principle of symmetry. Imposing hard constraints on GNN models via symmetry can facilitate the network convergence, enhance the model’s interpretability, and improve the ability of spatial feature representation learning.
Table S2. Basic information of datasets.
3D GS RD Cubic $[0,96]^{3}$ ($u,v$) 13,824 ($24^{3}$) 3,000 10 (5/3/2) Periodic No
B.3.2. Temporal strategy
Predicting the temporal evolution of dynamics can be cast as a Markov process. A simple approach to model such an evolution is resorted to numerical discretization and integration over time. Explicit time integration schemes are favorable and preferred in deep learning (e.g., MPNNs (Pfaff et al., 2021)) in the context of temporal forecasting (e.g., rollout prediction schemes), thanks to their straightforward implementation in a computational graph. However, if the time step between two evolved states is too big, it may cause instability issue. We are motivated to address this issue by proposing a learnable sub-stepping method. To this end, we view the message-passing process of the feedforward SymMPNN layers as the temporal propagation of dynamics at a finite number of sub-steps. Hence, we design a trainable temporal integrator to sequentially march the latent features of the SymMPNN layers so as to improve the model’s performance for long-range temporal prediction.
Appendix C Dataset and Baselines
We evaluate the performance of our proposed method and baseline models on three simulated datasets and one real-world weather data. For the generated datasets, we consider three different types of PDE systems, including 2D viscous Burgers equation, 3D Gray-Scott Reaction-Diffusion (GS RD) equation, and 2D unsteady incompressible Cylinder Flow (CF). Moreover, the real-world dataset is the atmospheric dynamics in the Black Sea (BS) field, where no explicit governing equations are considered. The detailed information of generating specific datasets can be found in Table S3. We compare our model with four representative baseline models, outlined as follows. And we make every effort to strike a balance between the model performance and the parameter count (e.g., typically aiming to keep it within the range of 500,000 to 600,000). The parameters of the relevant baseline models are shown from Table S11 to Table S13. Notably, all the hyperparameters are obtained from various ranges. For example, the learning rate is selected from the set $[10^{-1},10^{-2},10^{-3},10^{-4},10^{-5}]$.
C.1. Dataset Description
C.1.1. 2D Burgers Equation
Convection-diffusion systems emerge from the interplay of the diffusion and convection (advection) processes, describing scenarios where particles, energy, or other physical attributes are transported in time and space. One of the most well-known PDEs used to model such phenomena is the viscous Burgers equation, commonly found in simplified fluid mechanics, nonlinear acoustics, and gas dynamics, expressed as (with the source term omitted) $\partial\mathbf{u}/\partial t=\mathbf{D}\nabla^{2}{\mathbf{u}}-\mathbf{u}\cdot\nabla\mathbf{u}$. Here, $\mathbf{D}\in\mathbb{R}^{d\times d}$ denotes the diffusivity tensor, $\nabla^{2}{\mathbf{u}}$ characterizes the diffusion, and $\mathbf{u}\cdot\nabla\mathbf{u}$ represents the convection (or advection). In this example, we consider a 2D incompressible flow ($\mathbf{u}=[u,v]^{T}$) described by the Burgers equation, subjected to a randomized IC in a squared domain with periodic BCs.
C.1.2. 3D Reaction-Diffusion system
The reaction-diffusion (RD) process is an interesting phenomenon commonly seen in nature, e.g., the interaction among chemical or biological substances, the formation of Turing patterns in the skin of tropical fishes, etc. Such a process emphasizes the role of intrinsic material diffusion and localized reactions in shaping system dynamics in space and time. Mathematically, it can be represented by the following PDE: $\partial\mathbf{u}/\partial t=\mathbf{D}\nabla^{2}{\mathbf{u}}+\mathcal{R}(\mathbf{u})$, where $\mathbf{u}\in\mathbb{R}^{d}$ represents the concentration variables, $\mathbf{D}\in\mathbb{R}^{d\times d}$ the diffusion coefficient matrix, and $\mathcal{R}(\mathbf{u})$ the reaction term. Herein, we consider a GS RD equation in a 3D space (e.g., $\mathbf{x}\in\mathbb{R}^{3}$), which describes the reaction and diffusion of two chemical species ($\mathbf{u}=[u,v]^{T}$). The reaction equations for this system read $\mathcal{R}_{u}=-{u}{v}^{2}+\alpha(1-{u})$ and $\mathcal{R}_{v}={u}{v}^{2}-(\alpha+\beta){v}$, where $\alpha$ and $\beta$ are the coefficients prescribing the reaction process.
C.1.3. 2D Flow past a cylinder
We further consider a more challenging example (e.g., 2D fluid flow past a cylinder) to test the generalization ability of CiGNN over BCs. In practice, cylinder flow analysis is indispensable in scenarios such as fluid around bridge foundations, offshore drilling platform supports, underwater pipelines, and jetty piles. Here, the incompressible fluid flows into a channel (1.2 m $\times$ 0.41 m) from left to right and passes a circular cylinder located near the left boundary, as depicted in Figure 3c. The flow dynamics can be modeled by the Navier-Stokes (NS) equation given by
(S27a) $\frac{\partial\mathbf{u}}{\partial t}+\nabla\cdot(\mathbf{u}\otimes\mathbf{u})$ $=\frac{1}{Re}\nabla^{2}\mathbf{u}-\frac{1}{\rho}\nabla p,$
(S27b) $\nabla\cdot\mathbf{u}$ $=0,$
where $\mathbf{u}=[u,v]^{T}\in\mathbb{R}^{2}$ denotes the vector field of fluid velocity, $\rho$ the fluid density, and $\otimes$ the dyadic tensor product. Note that the pressure field $p\in\mathbb{R}^{1}$ constrains the flow and serves as a Lagrange multiplier to enforce the divergence-free condition of the fluid in Eq. S27b. The Reynolds number $Re$ dictates the complexity of the fluid, e.g., turbulence occurs when $Re\gg 1$. For the particular example considered herein, the value of $Re$ is inversely proportional to the cylinder size. We herein consider the location of the circular cylinder varies (resulting in the change of BCs) for the generation of different datasets (see Figure 3c), while the inflow velocity remains constant. Hence, $Re$ is no longer a constant. As the fluid moves around the cylinder, the reduced cross-sectional area leads to accelerated velocity and decreased pressure along its path.
Table S3. Computational parameters for datasets generation. $\delta x$ and $\delta t$ denote the spatial and time spacing, respectively.
Dataset PDE Parameters Size $\delta t(s)$ $\delta x$
2D Burgers Eq. S28 $\nu=0.01,\lambda,\gamma\sim\mathcal{N}(0,1),{c}\sim\mathcal{N}(-1,1)$ $1,001\times 2\times 50^{2}$ 0.001 0.02
2D CF $\nu=0.001,U_{mean}=1m/s$ $1,001\times 3\times[3,000,5,000]$ 0.01
3D GS RD Eq. S30 $D_{u}=0.2$, $D_{v}=0.1$, $\alpha=0.025$, $\beta=0.055$ $3,001\times 2\times 48^{3}$ 0.25 2
2D BS
C.1.4. 2D Hydrological Black Sea Dataset
Despite satisfactory performance, the CiGNN model is only tested on simulated datasets for systems governed by well-posed PDEs in the previous examples. However, in real-world problems, the field observation data often contains vast uncertainties induced by measurement noise, variant boundary and source conditions, etc., posing a grand challenge for data-driven modeling of complex spatiotemporal dynamics. For example, predicting precipitation in a coastal area can be substantially affected by unforeseeable extreme weather events. In light of the practical application of CiGNN, we have intentionally considered a real-world problem in the field of oceanography, namely, forecasting the hydrological dynamics of the Black Sea.
The Black Sea dataset consists of monthly and daily ocean fields of the BS basin, e.g., the daily averaged water flow velocities $u(x,y,t)$ and $v(x,y,t)$ and sea surface temperature ${T}(x,y,t)$, recorded from June 1, 1993 to June 31, 2021, with a horizontal resolution of $1/27^{\circ}\times 1/36^{\circ}$ and 31 levels of sea elevation111 https://data.marine.copernicus.eu/product/BLKSEA_MULTIYEAR_PHY_007_004/description. It is noted that the multi-year field observations located in an irregular domain encompass an intrinsic nature, e.g., low spatial and temporal resolution, limited long-term data sequence, inherent noise, data gaps, and unknown environmental influences such as immeasurable effects of climate change, among others, which would result in critical challenges in establishing a robust learning model for accurate forecasting. Nevertheless, considering the periodic patterns of the hydrological time series, we encode the cyclical timestamp into our CiGNN model to capture effectively the variations in specific portions of the dataset.
C.2. Dataset Generation
C.2.1. 2D Burgers Equation

### v2

@@ 5.2. Loss function and Evaluation metrics
where $cov[\cdot,\cdot]$ the covariance function; $\sigma_{\mathbf{u}}$ and $\sigma_{\hat{\mathbf{u}}}$ the standard deviations of $\mathbf{U}$ and $\widehat{\mathbf{U}}$, respectively. In addition, introducing noise to the training data can stabilize the model and improve the generalization ability for learning spatiotemporal dynamics (Sanchez-Gonzalez et al., 2020; Pfaff et al., 2021). Therefore, we add a small amount of Gaussian noise in the training data to improve the CiGNN’s performance.
5.3. Treatment of Boundary Condition
Figure 3. Error distributions and the system state snapshots predicted by CiGNN and other baselines. a, The 2D viscous Burgers equation example. b, The 3D Gray-Scott equation example. c, The 2D CF example setup with the fluid flowing in from the inflow to the outflow.
@@ 5.3. Treatment of Boundary Condition
5.4. Training settings and resources
For a fair comparison, we train each model ten times and select the one with the median performance according to the evaluation of the validation datasets. All models are trained with the Adam Optimizer (Kingma and Ba, 2015) and the ReduceLROnPlateau learning scheduler (Paszke et al., 2019) which decays the learning rate by a factor of $0.8$. We train all models for 1,000 epochs and set early stopping with a patience of 100 epochs. We utilize the GELU activation function (Paszke et al., 2019) for all MLP hidden layers, while linear activation is considered for the output layers. Additionally, layer normalization is applied after each MLP to improve the training convergence, except the Decoder.
We use the similar architecture as in (Pfaff et al., 2021) but only keep the encoder and decoder block. For concretely, the encoder module consists of a 2-layer MLP with a hidden size of 128, and the decoder has a 2-layer MLP with a hidden size of 128. We only use 4 SymMPNN blocks for the processor, where 2 layer MLPs with a hidden size of 128 are employed as the node, edge and flux update functions. The same hyperparameters are used in MGN, MP-PDE, and CiGNN. Notably, all the hyperparameters are selected from various ranges. For example, the learning rate is selected from the set $[10^{-1},10^{-2},10^{-3},10^{-4},10^{-5}]$. For hardware, the models were trained on the NVIDIA A100 GPU (80GB) on an Intel(R) Xeon(R) Platinum 8380 CPU (2.30GHz, 64 cores) server.
5.5. Results
[…]
Prediction results of CiGNN on real-world BS hydrological dataset.
Table 3. Ablation results on four datasets. Here, the symbol “–” represents that Task #$\#2$ on CF dataset does not exist since multi-mesh strategy is not utilized. The bold values and underlined values represent the optimal and sub-optimal results.
@@ 5.5. Results
Table 4. Summary of performance of CiGNN and its variants. CiGNN- and CiGNN are two variants of CiGNN, where in the SymMPNN block only the asymmetric information passing is considered.
5.5.1. 2D Convection-Diffusion system
The results listed in Table 1 show that CiGNN consistently outperforms other baseline models in the context of prediction accuracy. Figure 3a displays the distribution of prediction errors as well as the predicted snapshots at a typical time step (0.995 s) by different methods. The CiGNN model lifts the prediction accuracy by 1$\sim$2 orders of magnitude in terms of MSE metric compared with all other baselines. Due to the meshless and point-wise nature of DeepONet, it falls short in scalable modeling of higher dimensional spatiotemporal dynamics. Given the squared domain with structured grid mesh, FNO produces a moderate prediction result in the situation with limited training data. The performance of MP-PDE aligns with that of MGN. Note that MP-PDE is a derivative of MGN with the primary distinction in their decoder components, e.g., MGN employs multilayer perceptron (MLP) while MP-PDE utilizes convolution for decoding. Despite discrepant performance of other models, CiGNN generalizes well to various ICs and maintains a high prediction accuracy, which demonstrates notable strength and shows clear superiority over all baselines on 2D Burgers equation.
5.5.2. 3D Reaction-Diffusion system
5.5.3. 2D Flow past a cylinder
5.5.4. 2D Hydrological dynamics forecast of the Black Sea
@@ 5.6. Ablation study
To gain deeper insights into how each component of the model affects its overall performance, we conducted an ablation study based on the aforementioned datasets. The results in Table 3 show that the removal of each module leads to the performance degradation. We summarize the observations as follows:
@@
We also consider two variants of CiGNN, namely, CiGNN- and CiGNN, where only the asymmetric information passing is considered in the SymMPNN blocks. As depicted in Table 4, the result shows the efficacy of the symmetric information-passing component. In the case of BS dataset, CiGNN outperforms CiGNN, which indicates that the effective learning of the inequality term in real-world scenarios is not fully guaranteed.
Moreover, we provide the ablation results on comparison of various temporal integration strategies shown in Table 5. It is evident that a 4-layer CiGNN model with the simplest 1D convolution filter-based strategy (denoted by C3, with minimal parameter overhead) achieves the best performance, surpassing other strategies including the 2D convolution filtering, TCN and LSTM, demonstrating that our temporal strategy (1D Conv filter) plays a key role in CiGNN to improve the model’s long-term rollout prediction accuracy.
@@
Case No. of layers Temporal strategy No. of parameters RMSE $\downarrow$
@@ Acknowledgements.
The work is supported by the National Natural Science Foundation of China (No. 62276269, No. 92270118), the Beijing Natural Science Foundation (No. 1232009), and the Strategic Priority Research Program of the Chinese Academy of Sciences (No. XDB0620103). H. Sun and Y. Liu would like to thank the support from the Fundamental Research Funds for the Central Universities (No. 202230265 and No. E2EG2202X2). P. Ren would like to disclose that he was involved in this work when he was at Northeastern University, who has not been supported by Huawei Technologies.
Appendix A Conservative Property of CiGNN
@@
Theorem 1: If there is a continuously differentiable tensor field $\mathbf{F}$ on a neighborhood of the volume $\mathcal{V}$, then we have that a volume integral over the volume $\mathcal{V}$ is equal to the surface integral over the boundary of the volume $\mathcal{V}$:
@@
where $flux_{i}$ is the total flow quantity through the node $i$; $flow_{i\cdot}$ and $flow_{\cdot i}$ the flows that leave and enter the node $i$. Subsequently, an easy corollary of Theorem 1 is described as follows.
Corollary 1: If the volume ${\mathcal{V}}$ can be partitioned into some separate parts $\mathcal{V}_{1},\mathcal{V}_{2},\dots,\mathcal{V}_{n}$, the flux out $\Phi(\mathcal{V})$ of the original volume ${\mathcal{V}}$ should be equal to the sum of flux through each component volume:
@@
where the edge weight function $w_{ij}$ associates to every edge ${e}_{ij}$ a value. Similarly, $w_{ji}$ also associates to every edge ${e}_{ji}$ a value.
Definition 2: Based on the weighted graph derivative, we define the linear weighted gradient operator $\nabla_{w}:\mathcal{H}(V)\rightarrow\mathcal{H}(E)$ as:
@@
where $\mathcal{N}_{i}$ denotes the neighborhood node’s index set for node $i$. For simplicity, we simplified $F({u}_{i},{u}_{j})$ and $F({u}_{j},{u}_{i})$ as $F_{ij}$ and $F_{ji}$.
@@
where ${s}(\cdot)$ represents the source term. Then, substituting Eqs. S6 and S8 into Eq. S9, we get the following equation:
@@
We realize that the conservation term can be formulated by the skew-symmetric tensor and the inequality by the symmetric tensor (Tadmor, 1984). Instead of constructing such tensors directly, we implicitly build a learnable scheme to retain the above structure property:
@@
Appendix B Dataset Generation
B.1. 2D Burgers Equation